OpenStudy (anonymous):

Medal & fan (: How do the graphs of y = 1/x and y=5/x+6 compare?

OpenStudy (anonymous):

y = 5/x + 6 y = 5/x stretches out the graph by 5 times. It increases and decreases 5 times more. y = 5/x + 6 makes your raises your graph 6 units higher.

OpenStudy (anonymous):

@britneyspears alright thanks (:

OpenStudy (anonymous):

You're welcome!! Xo!